# Strategic Analysis of Fair Rank-Minimizing Mechanisms with Agent Refusal Option

### v3

@@ 1 Introduction
The rank distributions of the assignments are seriously considered by many real world matching authorities. For example, Featherstone (2020) mentions that Teach for America considers the rank distribution when selecting its assignment. Therefore, the rank-minimizing mechanisms are focused on by several recent studies such as Featherstone (2020), Nikzad (2022), Ortega and Klein (2023) and Troyan (2024). Some of them highlight several advantages of the mechanism with regard to the efficiency of assignment. Featherstone (2020) shows that any rank-minimizing mechanism satisfies another property of efficiency called ordinal efficiency, first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the result of a rank-minimizing mechanism and those of the DA mechanism and the TTC mechanism, and show that the result of a rank-minimizing mechanism improves upon those of others in several terms of efficiency.
@@
Next, we consider the case where agents can refuse their assignment and instead obtain the outside option. Featherstone (2020) firstly considers this situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option. In contrast, we assume that each agent has complete information on the capacities of all object types. For example, in a school choice market, information about all schools, including their capacities, is assumed to be disclosed to all students. Additionally, we do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents because we focus only on strategic dominance relations.
@@
Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy. We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair. That is, to ensure that any outside option demotion strategy can fail, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types. However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the (best) outside option of an agent. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
@@
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefers $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
Assignment $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ if each agent $a$ with $p_{a}$ weakly prefers $x$ to $x^{\prime}$. If $x$ weakly ordinally dominates assignment $x^{\prime}$ for $p$ and (1) is satisfied with inequality for at least one agent–rank pair, then $x$ ordinally dominates assignment $x^{\prime}$ for $p$. An assignment that is not ordinally dominated by any other assignments is said to be ordinally efficient for $p$.
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. Any wasteful assignment for $p$ must not be ordinally efficient. [^1] Moreover, a wasteful assignment for $p$ must not be efficient in a weaker efficiency concept such as ex-post efficiency, which is discussed several previous studies such as Bogomolnaia and Moulin (2001). Thus, not being wasteful is a very weak requirement for efficiency.
For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
which is the sum of expected rank. An assignment $x^{\ast}\in\mathcal{X}$ is rank efficient for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank efficient assignments for $p$.
The following result is shown by Featherstone (2020).
Remark 2
Any rank efficient assignment for $p$ is ordinally efficient for $p$.
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank efficient for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
@@
and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank efficient, then $y^{i}$ is also rank efficient for all $i=1,\cdots,I$.
@@
There must exist at least one deterministic rank efficient assignment.
Nikzad (2022) and Troyan (2024) consider a deterministic rank efficient assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank efficient in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
@@ 3 Rank-minimizing Mechanism
A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$.
@@
We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank efficient for all $p$ and satisfies ETE. The following mechanism is one of them.
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank efficient assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
@@
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank efficient for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
@@
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th best one.
@@
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the best type and the second best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
@@
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$. Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a fair rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.
@@ 4 Refusal option
In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game. In the third stage, all agents decide whether accept or refuse the object type that is assigned to them. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). In this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours. [^3] Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders.
@@
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their best object $o_{1}$. On the other hand, suppose
@@
for $a^{\prime}=a_{2},a_{3}$. Thus, it is ordinally dominated by
$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right).$
@@
By this result, an agent is never hurt by using an ODS instead of truth-telling. Next, we show that with some conditions, there exists some ODS such that an agent may be better off by using it instead of truth-telling.
@@
Theorems 1 and 2 do not say that an ODS is a dominant strategy. However, when if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
@@
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second best type increases it, when $o_{3}$ is the first best type of another agent. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
@@
if $\left|\mathcal{A}\right|>q_{o}+1$, then the first best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
@@
First, we explain why the strategic issues mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
@@
Let $R\left(\varnothing,r_{a}\right)=k$. Suppose $k\geq 3$; that is, $\varnothing$ is neither the best type nor the second best type for $r_{a}$. Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$. For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$. We assume

### v4

@@ 1 Introduction
The rank distributions of assignments are a critical consideration for many real-world matching authorities. For instance, Featherstone (2020) notes that Teach For America takes rank distribution into account when making assignment decisions. Consequently, rank-minimizing mechanisms are discussed in recent studies, including Featherstone (2020), Nikzad (2022), Ortega and Klein (2023), Feizi (2024), and Troyan (2024). Several of these studies highlight the advantages of rank-minimizing mechanisms, particularly regarding the efficiency of assignments. By the results of Featherstone (2020) and Feizi (2024), any rank-minimizing mechanism satisfies other efficiency properties, such as ordinal efficiency, which is first introduced by Bogomolnaia and Moulin (2001). Moreover, Ortega and Klein (2023) compare the outcomes of rank-minimizing mechanisms with those of the DA mechanism and the TTC mechanism, showing that rank-minimizing mechanisms improve upon the others in several aspects of efficiency.
@@
Next, as the main analysis, we consider the case where agents have the refusal options. [^1] Feigenbaum et al. (2020) and Afacan (2022) state that in some real school choice markets, such as public school markets in Boston, New York City, and Turkey, students actually refuse their assignments. For example, about 10% of the students withdrew from the New York City public high school placement. Featherstone (2020) theoretically considers the situation, but focuses only on the case where each agent has limited information on both the preferences of other agents and the capacities of all object types except the outside option. In contrast, we assume that each agent has complete information on the capacities of all object types. For example, we consider a school choice market where information about all schools, including their capacities, is disclosed to all students. We do not need any assumptions about the information structure of agents regarding the preferences and rationality of other agents, because we focus only on strategic dominance relations.
@@
Due to the negative result on the uniform rank-minimizing mechanism introduced above, we modify it so that truth-telling is not strategically dominated by any other strategy. We call this mechanism the modified uniform rank-minimizing mechanism, which is shown to be fair. That is, to ensure that any outside option demotion strategy can fail in some case, in the mechanism, increasing acceptable object types can make the agent less likely to be assigned their acceptable types. However, using the modified uniform rank-minimizing mechanism may create another strategic issue: an agent may have an incentive to reduce the number of truly acceptable object types.
@@ 2 Model
We consider the model discussed by Kojima and Manea (2010) and Featherstone (2020). Let $\mathcal{A}$ and $\mathcal{O}$ be finite sets of agents and object types respectively. We assume $\left|\mathcal{A}\right|\geq 2$ and $\left|\mathcal{O}\right|\geq 2$. Each agent must be allocated one copy of an object whose type is an element of a set denoted by $\mathcal{O}$. Let $q_{o}\geq 1$ be the capacity of an object type $o\in\mathcal{O}$. We assume that there is a null object type denoted by $\varnothing\in\mathcal{O}$ satisfying $q_{\varnothing}\geq\left|\mathcal{A}\right|$. That is, the null object type can be assigned to all agents simultaneously. The null object type $\varnothing\in\mathcal{O}$ can be interpreted as the first-best alternative object of an agent outside of this market. In Section 4, we consider the situation where the null object type has the other characteristic. To avoid trivial cases, we assume $q_{o}\in\left[1,\left|\mathcal{A}\right|\right)$ for all $o\in\mathcal{O\setminus}\left\{\varnothing\right\}$.
Let $\mathcal{P}$ be the set of all possible strict rankings (linear orders) of the elements in $\mathcal{O}$. Let $p_{a}=\left(p_{a}^{1},\cdots,p_{a}^{\left|\mathcal{O}\right|}\right)\in\mathcal{P}^{\left|\mathcal{A}\right|}$ be a preference order revealed by $a$, where $p_{a}^{k}=o$ means that type $o$ object is ranked $k$th-best one under $p_{a}$. For notational convenience, we let $R:\mathcal{O\times P}\rightarrow\left\{1,\cdots,\left|\mathcal{O}\right|\right\}$ be a rank function; that is, $p_{a}^{k}=o$ if and only if $R\left(o,p_{a}\right)=k$. Moreover, if $R\left(o,p_{a}\right)<R\left(\varnothing,p_{a}\right)$ (resp. $R\left(o,p_{a}\right)>R\left(\varnothing,p_{a}\right)$), then $o$ is said to be acceptable (resp. unacceptable) for $p_{a}$. Thus, $R\left(\varnothing,p_{a}\right)-1$ represents the number of acceptable object types for $p_{a}$. Let $r_{a}\in\mathcal{P}$ be the true preference order of agent $a$. Especially, if $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, then $o$ is said to be truly acceptable for $a$.
@@
for all $k=1,2,\cdots,\left|\mathcal{O}\right|$, then agent $a$ with $p_{a}$ is said to weakly prefer (in the sense of first-order stochastic dominance) $x$ to $x^{\prime}$. Moreover, $a$ is said to strictly prefer $x$ to $x^{\prime}$ if $a$ weakly prefers $x$ to $x^{\prime}$ and (1) is satisfied with inequality for some $k=1,2,\cdots,\left|\mathcal{O}\right|-1$.
Assignment $x$ is wasteful for $p$ if there are $o,o^{\prime}\in\mathcal{O}$ and $a\in\mathcal{A}$ such that $\sum\nolimits_{a^{\prime}\in\mathcal{A}}x_{a^{\prime}o}<q_{o}$, $x_{ao^{\prime}}>0$ and $R\left(o,p_{a}\right)<R\left(o^{\prime},p_{a}\right)$. For an assignment $x\in\mathcal{X}$ and a preference profile $p=\left(p_{a}\right)_{a\in\mathcal{A}}\in\mathcal{P}^{\left|\mathcal{A}\right|}$, let $RV\left(x,p\right)$ be such that
@@
which is the sum of expected rank. An assignment $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$ if $RV\left(x^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Let $\mathcal{X}^{\ast}\left(p\right)$ be the set of rank-minimizing assignments for $p$. If $x^{\ast}\in\mathcal{X}$ is rank-minimizing for $p$, then $x$ is not wasteful for $p$.
The efficiency property of ”rank-minimizing” is stronger than several efficiency properties, such as rank efficiency and ordinal efficiency, analyzed in this literature (see Figure 1 in Feizi (2024)). Among them, not being wasteful is a very weak property.
In the subsequent section, we focus on the mechanisms whose input is $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$ and output is rank-minimizing for any $p\in\mathcal{P}^{\left|\mathcal{A}\right|}$. Before that we introduce some technical results.
@@
and $\sigma^{i}\in\left[0,1\right]$ for all $i=1,\cdots,I$. If $x^{\ast}$ is rank-minimizing, then $y^{i}$ is also rank-minimizing for all $i=1,\cdots,I$.
@@
There must exist at least one deterministic rank-minimizing assignment.
Nikzad (2022) and Troyan (2024) consider a deterministic rank-minimizing assignment; that is, they focus on a deterministic assignment $y^{\ast}\in\mathcal{Y}$ such that $RV\left(y^{\ast},p\right)\leq RV\left(y,p\right)$ for all $y\in\mathcal{Y}$. However, by Corollary 1, such $y^{\ast}$ must also be rank-minimizing in our definition; that is, $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for all $x\in\mathcal{X}$. Moreover, this result allows us to show that the mechanism introduced later is well-defined.
@@ 3 Rank-minimizing Mechanism
A mechanism used in the second stage is represented by $f:\mathcal{P}^{\left|\mathcal{A}\right|}\rightarrow\mathcal{X}$. First, we consider an efficiency property on mechanisms. We call $f$ a rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$.
@@
Other fairness properties discussed in the literature include weak envy-freeness and the equal division lower bound. However, as shown by Feizi (2024), both of these properties are incompatible with rank-minimizing. Since ETE is weaker than these fairness properties, it is compatible with rank-minimizing and is therefore adopted in this study.
@@
We focus on the efficient and fair mechanisms; that is, we call $f$ a fair rank-minimizing mechanism if $f(p)$ is rank-minimizing for all $p$ and satisfies ETE. The following mechanism is one of them.
Let $\mathcal{Y}^{\ast}\left(p\right)\subseteq\mathcal{Y}$ be the set of deterministic rank-minimizing assignments; that is, $y^{\ast}\in\mathcal{Y}^{\ast}\left(p\right),$ $RV\left(y^{\ast},p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$. By Corollary 1, $\mathcal{Y}^{\ast}\left(p\right)\neq\emptyset$. Let
@@
Thus, $RV\left(f^{U}\left(p\right),p\right)\leq RV\left(x,p\right)$ for any $x\in\mathcal{X}$; that is, $f^{U}\left(p\right)$ must be rank-minimizing for $p$. Since Nikzad (2022), Ortega and Klein (2023) and Troyan (2024) also discuss the uniform rank-minimizing mechanism, this is one of the most common rank-minimizing mechanisms in the literature.
@@
is satisfied for that $k$. In other word, we consider the case where the number of object types that are more preferable than the $k-1$th best one in $p_{a}$ is insufficient to distribute to all agents. In this case, for some $p_{-a}\in\mathcal{P}^{\left|\mathcal{A}\right|-1}$ and some fair rank-minimizing mechanism, $a$ is assigned to an object type that is less preferable or equal to the $k$th-best one.
@@
if $\left(p_{a_{1}},p_{a_{2}},p_{a_{3}}\right)$ is neither $\left(\psi^{\prime},\psi,\psi\right)$ nor $\left(\psi,\psi^{\prime},\psi\right)$ nor $\left(\psi,\psi,\psi^{\prime}\right)$. That is, if the first-best type and the second-best type of the three agents are respectively $o_{1}$ and $o_{2}$ and the third-best type of the two of them is $\varnothing$, then the agent revealing $o_{3}$ as the third-best one obtains $o_{1}$ with probability 00. Trivially, $f$ satisfies weak ETE. However, whether $f$ satisfies ETE or not is dependent on $q_{o_{2}}$.
@@
Second, suppose $q_{o_{2}}\geq 2$. In this case, $\bar{k}\left(\psi\right)=\bar{k}\left(\psi^{\prime}\right)=2$. Then, $a_{1}$ must have either $o_{1}$ or $o_{2}$ if it chooses either $\psi$ or $\psi^{\prime}$ and thus, the unique interest of $a_{1}$ is increasing the probability to obtain $o_{1}$. Thus, in this case, $r_{a_{1}}=\psi^{\prime}$ is strategically dominated by $\psi$. Therefore, even if a rank-minimizing mechanism is used, truth-telling may be strategically dominated by some other strategy. Moreover, in this case, $f$ does not satisfy ETE. First, since $q_{o_{1}}+q_{o_{2}}\geq\left|\mathcal{A}\right|$, any agent is assigned to neither $o_{3}$ nor $\varnothing$,∅ , as long as the assignment is not wasteful. Thus, $\psi$ and $\psi^{\prime}$ are essentially the same in this case. However, $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{1}}=0,$ but $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{2}}>0$ or $\left(f\left(\psi^{\prime},\psi,\psi\right)\right)_{a_{1}o_{3}}>0$ must be satisfied. Therefore, in this case, $f$ does not satisfy ETE.
@@ 4 Refusal option
In this section, we consider the following three-stage game, which is a modification of that in Section 3; that is, the first and second stages in this game are the same as in the previous game. In the third stage, all agents decide whether to accept or refuse the assigned object type. If an agent refuses their assignment, then the agent instead obtains the null object type $\varnothing$ (the outside option). That is, in this game, the null object type has two characteristics: first, there are sufficiently many copies of this type to allocate to all individuals, and second, agents can always obtain it at no cost as long as they abandon their assignment.
This third stage is also considered by Featherstone (2020) but is not explicitly addressed by almost previous studies related to ours. [^3] As exceptions, Feigenbaum et al. (2020) and Afacan (2022) consider an assignment problem where, after the initial allocation is decided, if someone refuses the assigned object, it is reallocated. Moreover, Doğan and Yenmez (2019) also explicitly consider this stage. However, this stage is implicitly considered by many of them; that is, they consider an ex-post individually rational (or more strictly an ex-post efficient) matching as a requirement and if an assignment is ex-post individually rational, no agent refuses their assignment. See, for example, Bogomolnaia and Moulin (2001) on the property.
@@
Featherstone (2020) defines the full extension of $r_{a}$denoted by $e_{a}$ in this study such that (i) $R\left(o,r_{a}\right)=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)<R\left(\varnothing,r_{a}\right)$, and (ii) $R\left(\varnothing,e_{a}\right)=\left|\mathcal{O}\right|$, and $R\left(o,r_{a}\right)-1=R\left(o,e_{a}\right)$ for all $o\in\mathcal{O}$ such that $R\left(o,r_{a}\right)>R\left(\varnothing,r_{a}\right)$; that is, $e_{a}\in\mathcal{D}_{a}$ is such that all non-null object types are placed in their true preference orders. Revealing the full extension of $r_{a}$ is one of the ODSs of $a$.
@@
Thus, revealing $d_{a_{1}}$ instead of $r_{a_{1}}$ may increase the probability that $a_{1}$ is assigned to their first-best object $o_{1}$. On the other hand, suppose
@@
for $a^{\prime}=a_{2},a_{3}$. Thus, if
$\tilde{x}=\left(\begin{array}[]{ccc}1/3&0&2/3\\ 1/3&1/2&1/6\\ 1/3&1/2&1/6\end{array}\right),$
then $a_{2}$ and $a_{3}$ strictly prefer and $a_{1}$ weakly prefers $\tilde{x}$ to $g\left(f^{U}\left(d_{a_{1}},\hat{p}_{a_{2}},\hat{p}_{a_{3}}\right),r\right)$.
@@
By this result, an agent is never hurt by using an ODS instead of truth-telling. Next, we show that with some conditions, there is some ODS such that an agent may be better off by using it instead of truth-telling.
@@
Theorems 1 and 2 do not imply that an ODS is a dominant strategy. However, if there is $a\in\mathcal{A}$ whose true preference satisfies (3) for some $o,o^{\prime}\in\mathcal{O}$, since truth-telling is strategically dominated by another strategy, $a$ has a strong incentive to manipulate its preference ranking, while one of the ODSs may not necessarily be chosen.
@@
We explain why $d_{a_{1}}$ strategically dominates $e_{a_{1}}$. In this example, $q_{o_{1}}+q_{o_{3}}=2<3=q_{o_{1}}+q_{o_{2}}$. That is, the copies of $o_{1}$ and $o_{2}$ are sufficient to allocate all agents, but those of $o_{1}$ and $o_{3}$ are insufficient. Thus, revealing $o_{2}$ as a second-best type does not increase the probability that $a_{1}$ is assigned to $o_{1}$. On the other hand, revealing $o_{3}$ as a second-best type increases it, when $o_{3}$ is the first-best type of another agent. Thus, revealing the full extension may be strategically dominated by another ODS when an unacceptable type (in this case $o_{3}$) is less preferable than the other (in this case $o_{2}$) but the number of copies of the former is less than that of the latter.
@@
if $\left|\mathcal{A}\right|>q_{o}+1$, then the first-best type of the other agents ($\left|\mathcal{A}\right|-\left(q_{o}+1\right)$ agents) represented by $a^{\prime\prime}$ is $\varnothing$, then
@@
First, we explain why the strategic issue mentioned in Example 2 does not arise in the modified uniform rank-minimizing mechanism. If $r_{a}=p_{a^{\prime}}$ where $p_{a^{\prime}}$ is defined in (ii), then $\left(d_{a},p_{-a}\right)$ satisfies (i), (ii) and (iii) where $d_{a}\in\mathcal{D}_{a}$. This implies that for some $p_{-a}$, if agent $a$ adopts an ODS instead of truth-telling and $f^{M}$ is used, then $a$ is worse off.
@@
Let $R\left(\varnothing,r_{a}\right)=k$. Suppose $k\geq 3$; that is, $\varnothing$ is neither the first-best type nor the second-best type for $r_{a}$. Thus, we let $r_{a}^{1}=o$ and $r_{a}^{2}=o^{\prime}$. For $p_{-a}$, there are $q_{o}$ agents represented by $a^{\prime}$ whose revealed preference order is such that $p_{a^{\prime}}^{k^{\prime}}=r_{a^{\prime}}^{k^{\prime}}$ for all $k^{\prime}=1,\cdots,k-2$ and $p_{a^{\prime}}^{k-1}=\varnothing$. We assume